aggyalysha:

Eloise started to solve a radical equation in this way: square root of negative 2x plus 1 − 3 = x square root of negative 2x plus 1 − 3 + 3 = x + 3 square root of negative 2x plus 1 = x + 3 square root of negative 2x plus 1 − 1 = x + 3 − 1 square root of negative 2 x = x + 2 (square root of negative 2 x)2 = (x − 4)2 −2x = x2 − 8x + 16 −2x + 2x = x2 + 8x + 16 + 2x 0 = x2 + 10x + 16 0 = (x + 2)(x + 8) x + 2 = 0 x + 8 = 0 x + 2 − 2 = 0 − 2 x + 8 − 8 = 0 − 8 x = −2 x = −8 Both solutions are extraneous because they don't satisfy the original equation. What error did Eloise make? She added 2x after squaring both sides. She subtracted 1 before squaring both sides. She factored x2 + 10x + 16 incorrectly. She did not check for extraneous solutions.

minustempo:

If I understand this correctly, the original equation is sqrt(2x+1) - 3 = x then it became sqrt(2x+1) = x+3 so the problem is that sqrt(2x+1) is not the same as sqrt(2x) + 1 so you can't subtract 1 without squaring both sides first

aggyalysha:

thanks i appreciate it :)